Assuming that the rest radius of earth is 6400 km and its orbital speed about the sun is 30 km/sec,


how much does earths diameter appear to be shortened to an observer on the sun, due to earth’s


orbital motion?



1
Expert's answer
2021-12-29T12:43:01-0500

Answer

Radius of earth 6400Km

Speed v=30km/sec

"β =\frac{ v}{c}\\

 = \frac{30 km s^{−1}}{3 × 10^5km^{ s−1 }}\\

= 10^{−4}"


Now

"\frac{1}{γ} = (1 − β^2)^{1/2 }\\

= (1 − \frac{1}{2} × β^2)"

So earth shrunk

"ΔL = L_0 − L \\

= L_0 −L_0/γ \\

= L_0 − L_0(1 − β^2)^{\frac{1}{2} }\\

= \frac{1}{2}L_0β^2 \\

= \frac{1}{2} ×6,400× 10^{−8} km\\

 = 3.2 cm"

Thus the earth appears to be shrunk by 3.2 cm